Write y = – x + 7 in standard form using integers.
how would you do this problem I 'am in 7th grade math

Answer by jim_thompson5910(35256)

Add x to both sides to get

y = – x + 7

y + x= – x + 7 + x

x + y = 7

So the answer is x + y = 7